# ViLLa: Video Reasoning Segmentation with Large Language Model

### v2

@@ 4.1 Model Design
Context Aggregation Module. In this module, we aim to aggregate text-related visual features and inject them to generate text embedding that could represent the current frame. With the text embeddings $\mathbf{x}_{txt}$ and visual features, we formulate the context-based aggregation as:
@@
where ‘CrossAttn’ refers to cross-attention operation, and $\mathbf{x}_{txt}$ is the query, while $F_{t}^{L}$ is the value and key. By conducting the cross-attention, we further incorporate the visual cues into the refined text embeddings $\textbf{E}_{t}=\text{Concat}\{\textbf{E}_{t}^{i}\}_{i=1}^{M}$. However, we argue that not all queries are needed for the refined text embeddings. Unlike QFormer which adopts 32 input queries as input LLM tokens, we propose to condense these embeddings into more concentrated tokens. After getting $\textbf{E}_{t}$ from the context aggregation between text embeddings and visual features, we select the output embeddings with $K$ higher-response scores from the attention matrix. That is, the final condensed input embeddings $\textbf{E}_{t}^{c}$ preserve the most relevant visual information related to user input.
Multi-level Segmentation Tokens. Aiming to enrich the encoding of target-specific information and thereby facilitate the generation of high-quality mask tracklets across temporal dimensions, we devise multi-scale segmentation tokens, representing both frame-level and video-level concepts, tailored to meet the characteristics of video segmentation tasks that demand the modeling of multiple target movements. Formally, we define $C_{\text{seg}}=\left\{c_{n}^{s}\in\mathbb{R}^{d}\right\}_{n=1}^{N}$, where $s\in\{f,v\}$ indicates the frame- or video-scale tokens, $N$ denotes the number of tokens per scale, and $d$ refers to the hidden dimension of LMMs. The multi-level segmentation tokens $C_{\text{seg}}$, combined with the visual features $F_{t}^{L}$, and condensed text embeddings $\textbf{E}_{t}^{c}$, are then processed by the LLM to autoregressively generate the response $y_{\text{out}}$:
@@
Using an example can provide a more vivid explanation of how Large Language Models generate responses. When the user inputs “segment the leftmost zebra in the video”, the output will include not only a textual answer but also the segmentation tokens. $C_{\text{seg}}$: “The zebra is $c_{1}^{f},c_{1}^{v}$” (considering the easiest case where $N=1$ for $C_{\text{seg}}$). Before the decoder, an additional projection layer $\Phi$ is utilized to adjust the dimensions of segmentation tokens, in the form of $Q^{s}=\Phi(C_{\text{seg}})$, ($s\in\{f,v\}$). The output multi-scale segmentation embeddings $Q^{v},Q^{f}$ derived from $C_{\text{seg}}$ are inputs to the pixel decoder $\mathcal{D}$ alongside visual features $F_{{t}}$ for mask tracklet generation.
Video-Frame Decoder. We design a novel Video-Frame decoder $\mathcal{D}$ to generate mask tracklets from the multi-scale visual features $F_{{t}}$ and segmentation tokens $Q^{v},Q^{f}$. Based on the design of the transformer decoder in Mask2Former, $\mathcal{D}$ consists of $L$ decoder layers, where $l^{\text{th}}$ layer cascades a masked cross-attention $h_{\text{CA}}^{l}$, a self-attention $h_{\text{SA}}^{l}$, and a feed-forward network $\text{FFN}^{l}$. Now, we have the two scales of segmentation embeddings, $Q^{f}$ focusing on every frame separately, while $Q^{v}$ interacting with the whole video features.
Thus both the frame- and video-level segmentation embeddings pass the transformer decoder. For each step, these two embeddings are updated as follows:
@@
$\mathbf{Q}^{l+1,v^{{}^{\prime}}}=\gamma\cdot\text{Softmax}(({Q}^{l+1,v}\times{{Q}^{l+1,f}}^{\top})\times{{Q}^{l+1,f}})+(1-\gamma)\cdot\mathbf{Q}^{l+1,v},$ (5)
@@
where $\mathbf{Q}^{l+1,v^{{}^{\prime}}}$ is the modulated video-level segmentation embeddings at $l^{\text{th}}$ layer. In this aggregation process, we aggregate the frame-level embeddings with higher responses to our video-level segmentation embeddings in a momentum-based manner. The momentum factor $\gamma$, which is empirically set to 0.03, governs the update of the video-level embeddings. This choice is based on the assumption that the aggregation of frame-level embeddings should not induce substantial changes to the overall video-level representations.
Training Objectives. The model is trained end-to-end using the text generation loss $\mathcal{L}_{txt}$ and the segmentation mask loss $\mathcal{L}_{mask}$. The overall objective $\mathcal{L}$ is the weighted sum of these losses, determined by $\lambda_{txt}$ and $\lambda_{mask}$:
$\mathcal{L}=\lambda_{txt}\mathcal{L}_{txt}+\lambda_{mask}\mathcal{L}_{mask}.$ (6)
@@
where mask loss $\lambda_{mask}$ can be further divided into binary cross-entropy loss and the dice loss for both video-level and frame-level: $\mathcal{L}_{mask}^{s}=\lambda_{ce}^{s}\mathcal{L}_{ce}^{s}+\lambda_{dice}\mathcal{L}_{dice}$ ($s\in\{f,v\}$), and $\mathcal{L}_{txt}$ is the auto-regressive cross-entropy loss for text generation.
@@ 5 Experiment
In this section, we first present the implementation details and the evaluation benchmarks as well as baselines, and then show the comparison results on different video perception and our proposed reasoning benchmarks. Finally, we ablate on the key components in ViLLa.
@@ 5.1 Implementation Details
We use pre-trained multimodal model from InternVideo2-1B wang2024internvideo2 and InternVideo2-6B, with LoRA adopted for efficient fine-tuning. For the visual encoder and LLM, we apply pre-trained UMT-L from the InternVideo2 stage1 model as the vision encoder and Vicuna-7B v0. Following BLIP2, we deploy QFormer using the pre-trained $BERT_{base}$. To generate multi-scale visual features, we adopt the ViT-Adapter to generate the necessary multi-scale features for the video-frame decoder. It is important to note that video segmentation differs from image segmentation in that video segmentation involves capturing inter-frame relationships across multiple frames during training. This added complexity results in higher GPU memory requirements compared to image segmentation. To address this challenge, an efficient version of the VIT-Adapter is applied, which helps to reduce the memory consumption of the video segmentation model by removing all injectors.
@@
We use 8 A100 GPUs for training 50 epochs (approximately 1.5 days). We use the AdamW optimizer with the learning rate and weight decay set to $52e-$ 5 and $0.02$, respectively. The batch size per device is set to 4, the input frame number is 8, the input resolution is 224, and the gradient accumulation step is set to 10. Further details are presented in appendix.
5.2 Benchmarks and Baselines
@@
Benchmarks. We evaluate ViLLA on benchmarks with video segmentation and question-answering evaluations: VideoReasonSeg, referring video segmentation (Refer-Youtube-VOS, Ref-DAVIS17), conventional video segmentation (Youtube-VIS series), and VQA datasets in supplementary files (FunQA, Perception Test, CLEVRER). Through this evaluation, we validate the versatility of ViLLa in diverse segmentation tasks and question-answering tasks. For the VideoReasonSeg benchmark, we use the average precision (AP) defined in video segmentation as well as multiple choices to evaluate the video reasoning capabilities of models. In video segmentation, we formulate the queries following the dataset’s annotations in the format of “Please segment the <description>”, with <description> corresponding to the descriptions of the instance categories. In referring video segmentation, we follow previous methods to measure the $\mathcal{J}\&\mathcal{F}$. Training involves random sampling from VideoReasonSeg, VQA data, Video instance segmentation, and Referring video segmentation datasets. The model also follows the training setting of PixelLM ren2023pixellm to pre-train on image datasets to provide an initial value for segmentation tokens.
Baselines. To our best knowledge, ViLLa is the first to handle video reasoning segmentation. To demonstrate the effectiveness of ViLLa, we establish strong baselines for comparative analysis on the mentioned benchmarks. These baselines are developed from previous reasoning segmentation approaches (LISA, PixelLM), referring video segmentation methods (ReferFormer, MTTR), and video segmentation models (VITA, Mask2Former).
Table 1: Reasoning video segmentation results among ViLLa and previous related works on VideoReasonSeg benchmark. “Segm” refers to ‘Segmentation’ while “MC” indicates “Multiple Choices”.
Method Val Test
Segm MC Segm MC
AP AR Accuracy AP AR Accuracy
Mask2Former cheng2022masked 37.9 33.2 35.8 36.5 32.8 34.3
VITA heo2022vita 39.3 34.9 36.5 38.8 34.1 35.7
MTTR botach2022end 38.3 32.9 36.4 36.9 32.7 35.0
ReferFormer wu2022language 39.5 34.4 37.0 39.0 33.9 35.9
LISA-Llama-7B lai2023lisa 18.3 14.9 34.5 16.7 12.9 33.0
LISA-Llama2-13B lai2023lisa 20.1 15.4 35.2 18.5 13.6 34.7
PixelLM-Llama2-7B ren2023pixellm 33.2 29.5 42.5 30.4 29.7 37.2
PixelLM-Llama2-13B ren2023pixellm 34.2 30.1 44.6 31.8 29.9 41.3
ViLLa-InternVideo2-1B wang2024internvideo2 42.5 35.5 50.1 39.8 37.4 47.8
ViLLa-InternVideo2-6B wang2024internvideo2 44.3 39.2 53.4 41.2 38.8 50.5
@@ 5.3 Main Results
Video Reasoning Segmentation. Tab. 1 compares ViLLa with adapted competing methods on our video reasoning segmentation task. For the image reasoning segmentation method, to adapt to the video segmentation scenario, we add an extra tracker based on SAM-Track cheng2023segment . Since the original LISA only has one segmentation token, we calculate the AP between the predicted one tracklet and the ground truth tracklets. As for the VIS methods, we add the LLM to maintain their capability of complex reasoning.
@@
As shown in Tab. 1, our method outperforms the previous image reasoning segmentation approaches by a very large margin, demonstrating that our model is capable of accomplishing tasks involving complex reasoning across time. Because video reasoning, different from referring video segmentation and image reasoning segmentation, requires the model to possess reasoning ability based on the whole video. Thus, without comprehensive knowledge of the whole video, the model can not perform well. Also, we compared our method with two-stage methods which use LLM to generate text while using a video segmentation model to produce the tracklet. The results also indicate that our method exceeds these two-stage methods significantly. We argue that this is because we are trained end-to-end, while these methods decouple the video modeling and reasoning. Also, our model improves the association between video and frame via our specially designed context aggregation strategy and decoder structures, which distinguishes our method from previous works.
Table 2: Comparison on Ref-YouTube-VOS and Ref-DAVIS17.
Method Ref-YouTube-VOS Ref-DAVIS17

### v3

@@ 3.3 Context Synthesizer
This module aims to aggregate text-related visual features and inject them to generate text embedding representing the current frame. With the text embeddings $\mathbf{x}_{txt}$ and visual features, context-based aggregation is formulated as:
@@
where ‘CrossAttn’ refers to cross-attention operation, and $\mathbf{x}_{txt}$ is the query, while $F_{t}^{L}$ is the value and key. By the cross-attention, we further synthesize the visual cues into the refined text embeddings $\textbf{E}_{t}$. However, we argue that not all queries are needed for the refined text embeddings. Unlike QFormer which adopts 32 queries as input LLM tokens, we propose to condense these embeddings into more compact ones. After getting $\textbf{E}_{t}$, we select the output embeddings with $K$ highest-response scores from the attention matrix out of $M$. Therefore, the final input embeddings $\textbf{E}_{t}^{c}$ preserve the most relevant visuals.
Multi-level Segmentation Tokens. To enrich the target-specific encoding and thereby facilitate the generation of high-quality mask tracklets across frames, we devise multi-scale segmentation tokens, representing both frame-level and video-level concepts, tailored to meet the characteristics of video segmentation tasks that demand the modeling of multiple target movements. Formally, we define $C_{\text{seg}}=\left\{c_{n}^{s}\in\mathbb{R}^{d}\right\}_{n=1}^{N}$, where $s\in\{f,v\}$ indicates the frame- or video-scale tokens, $N$ denotes the number of tokens per scale, and $d$ refers to the hidden dimension of LMMs. The multi-level segmentation tokens $C_{\text{seg}}$, combined with the visual features $F_{t}^{L}$, and condensed text embeddings $\textbf{E}_{t}^{c}$, are then processed by the LLM to generate the response $y_{\text{out}}$:
@@
Using an example can provide a more vivid explanation of how Large Language Models generate responses. When the user inputs “segment the leftmost zebra in the video”, the output will include not only a textual answer but also the segmentation tokens. $C_{\text{seg}}$: “The zebra is $c_{1}^{f},c_{1}^{v}$” (considering the easiest case where $N=1$ for $C_{\text{seg}}$). Before the decoder, an additional projection layer $\Phi$ is utilized to adjust the dimensions of segmentation tokens, in the form of $Q^{s}=\Phi(C_{\text{seg}})$, ($s\in\{f,v\}$). The output multi-scale segmentation embeddings $Q^{v},Q^{f}$ derived from $C_{\text{seg}}$ are inputs to the pixel decoder $\mathcal{D}$ with $F_{{t}}$ for mask tracklet generation.
3.4 Hierarchical Temporal Synchronizer
We design a novel hierarchical temporal synchronizer $\mathcal{D}$ to generate mask tracklets from the multi-scale visual features $F_{{t}}$ and segmentation tokens $Q^{v},Q^{f}$. Derived from the design of the transformer decoder in Mask2Former, $\mathcal{D}$ consists of $L$ decoder layers, where $l^{\text{th}}$ layer cascades a masked cross-attention $h_{\text{CA}}^{l}$, a self-attention $h_{\text{SA}}^{l}$, and a feed-forward network $\text{FFN}^{l}$. Now, we have two scales of segmentation embeddings, $Q^{f}$ focusing on every frame separately, while $Q^{v}$ interacting with the whole video features. Both the frame- and video-level segmentation embeddings pass the transformer decoder, and are updated as follows:
@@ 3.4 Hierarchical Temporal Synchronizer
$\mathbf{Q}^{l+1,v^{\prime}}$ $=\gamma\cdot\text{Softmax}\left((\mathbf{Q}^{l+1,v}\times\mathbf{Q}^{l+1,f\top})\times\mathbf{Q}^{l+1,f}\right)$
$\quad+(1-\gamma)\cdot\mathbf{Q}^{l+1,v}$ (5)
@@
where $\mathbf{Q}^{l+1,v^{{}^{\prime}}}$ is the modulated video-level segmentation embeddings at $l^{\text{th}}$ layer. In this aggregation process, we aggregate the frame-level embeddings with higher responses to our video-level segmentation embeddings in a momentum-based manner. The momentum factor $\gamma$, empirically set to 0.03, governs the update of the video-level embeddings. This choice is based on the assumption that the aggregation of frame-level embeddings should not induce substantial changes to the overall video-level representations.
@@
Training Objectives. The model is trained end-to-end using text generation loss $\mathcal{L}_{txt}$ and segmentation mask loss $\mathcal{L}_{mask}$. The overall objective $\mathcal{L}$ is weighted sum of these losses, determined by $\lambda_{txt}$ and $\lambda_{mask}$: $\mathcal{L}=\lambda_{txt}\mathcal{L}_{txt}+\lambda_{mask}\mathcal{L}_{mask}$, where mask loss $\lambda_{mask}$ is divided into binary cross-entropy loss and dice loss for both video-level and frame-level: $\mathcal{L}_{mask}^{s}=\lambda_{ce}^{s}\mathcal{L}_{ce}^{s}+\lambda_{dice}\mathcal{L}_{dice}$ ($s\in\{f,v\}$), and $\mathcal{L}_{txt}$ is auto-regressive cross-entropy loss for text generation.
Table 1: Video Reasoning Segmentation on VideoReasonSeg benchmark. “Seg” refers to “Segmentation” while “MC” indicates “Multiple Choices”. ViLLa consistently outperforms other previous methods with different backbones and on different metrics by a large margin. Noted that previous methods are retrained to include our new proposed VideoReasonSeg in the training set.
Method Short Long Overall
Seg MC Seg MC Seg MC
$\mathcal{J}$$&\mathcal{F}$ Accuracy $\mathcal{J}$$&\mathcal{F}$ Accuracy $\mathcal{J}$$&\mathcal{F}$ Accuracy
MTTR [5] 17.2 - 12.7 - 15.6 -
ReferFormer [65] 17.5 - 13.4 - 16.0 -
LMPM [13] 26.4 - 18.5 - 23.1 -
OnlineRefer [63] 26.9 - 18.7 - 16.3 -
LISA-LLaVA-7B [24] 43.8 34.5 27.3 29.5 38.4 32.1
LISA-LLaVA-13B [24] 45.0 35.2 28.1 30.7 39.2 32.9
VideoLISA*-InternVideo2-1B [68] 51.0 38.4 32.4 32.9 45.3 36.7
VideoLISA*-InternVideo2-6B [68] 54.8 43.6 33.8 34.9 47.8 39.2
VISA-LLaVA-7B [68] 48.9 37.5 30.4 31.9 43.1 35.2
VISA-LLaVA-13B [68] 54.6 40.1 33.8 33.2 47.4 37.9
VISA-Chat-UniVi-13B [68] 54.8 42.6 34.0 34.3 47.6 39.6
VISA-InternVideo2-6B [68] 57.9 44.6 36.2 35.8 50.5 40.5
ViLLa-InternVideo2-1B [60] 57.4 46.5 38.9 37.6 51.2 43.1
ViLLa-InternVideo2-6B [60] 62.6 53.8 42.0 41.0 55.4 49.9
Table 2: Referring Video Object Segmentation on Ref-YouTube-VOS, Ref-DAVIS17, and MeViS. ViLLa consistently outperforms other methods on different benchmarks and on different metrics by a large margin.
Method Ref-YouTube-VOS Ref-DAVIS17 MeViS
[…]
URVOS [48] 47.2 45.3 49.2 51.6 47.3 56.0 27.8 25.7 29.9
LBDT [14] 49.4 48.2 50.6 54.1 - - 29.3 27.8 30.8
ReferFormer [65] 62.9 61.3 64.6 61.1 58.1 64.1 31.0 29.8 32.2
LMPM [13] - - - - - - 37.2 34.2 40.2
OnlineRefer [63] 62.9 61.0 64.7 62.4 59.1 65.6 - - -
DsHmp [19] 67.1 65.0 69.1 64.9 61.7 68.1 46.4 43.0 49.8
LISA [24] 54.4 54.0 54.8 66.0 63.2 68.8 37.9 35.8 40.0
PixelLM [46] 55.0 54.3 55.7 66.7 63.4 70.0 38.7 36.3 41.1
VideoLISA [4] 61.7 60.2 63.3 67.7 63.8 71.5 42.3 39.4 45.2
VISA [68] 63.0 61.4 64.7 70.4 67.0 73.8 44.5 41.8 47.1
ViLLa 67.5 64.6 70.4 74.3 70.6 78.0 49.4 46.5 52.3
Table 3: Video Reasoning Segmentation on ReVOS.
Method ReVOS
[…]
LISA-LLAVA-13B 41.8 39.6 43.9
VISA-Chat-UniVi-13B 50.9 48.8 52.9
VISA-InternVideo2-6B 52.4 50.1 54.7
ViLLa-InternVideo2-6B 57.0 54.9 59.1
3.5 VideoReasonSeg
For a more comprehensive evaluation of VRS, we establish VideoReasonSeg, a dataset containing text instructions, corresponding high-quality masks as well as multiple choices. To ensure reliable assessment, we collect a diverse set of videos from Youtube-VIS [70], OVIS [43], LV-VIS [56], MovieChat [50], and VideoMME [16]. Videos are categorized based on their duration, including short ($<1$ minutes) and long ($>5$ minutes). To fully evaluate the reasoning capabilities, we design two types of evaluation standards: 1) multiple-choice QAs; 2) instructions and answers. Considering that not all the annotations of selected datasets follow the multiple-choice QA format, we automatically convert the video annotations into this format via LLMs. This is mainly because the open-ended answers have to be scored by LLMs or user studies, which may either introduce evaluation bias or manual intervention. Ultimately, we produce 2 multiple-choice QA pairs for each of the video clips. As for the answers, we mainly focus on the quality of segmentation masks. The resulting VideoReasonSeg benchmark comprises a total of 3k videos with 15k object-instruction pairs. All videos are partitioned into two splits: train and val, containing 2k (1820 short videos and 180 long ones) and 1k videos (880 short videos and 120 long ones) respectively. The dataset incorporates three distinct complex scenarios: multiple object segmentation ($\sim$800), rapid motion ($\sim$200), and occlusion in long videos ($\sim$300). The details of data annotation are given in the appendix.
Evaluation Metrics. We follow previous works  [48, 13] to adopt $\mathcal{J}\&\mathcal{F}$ as the main evaluation metric, which is the average of region similarity $\mathcal{J}$ and contour accuracy $\mathcal{F}$. Besides, we adopt accuracy as the metric for multiple choices.
Dataset Generation Pipeline. We utilize GPT-4V for generating questions and multiple choices. Specifically, we feed all the instance category names and corresponding masks in the video to GPT-4V. With our carefully designed prompts, GPT-4V autonomously selects instances to construct question-answer pairs based on the video contents. Examples of such prompts are illustrated in the appendix.
@@ 4 Experiment
In this section, we first present the implementation details and the evaluation benchmark, and then show the comparison results on our proposed reasoning benchmarks and referring segmentation benchmarks. Finally, we ablate on the key components in ViLLa.
Table 4: Vision backbones. InternVideo2 demonstrates better performance (1.5 and 2.4 increase) compared to previous backbones.
Vision Backbone Segmentation Multiple Choices
$\mathcal{J}$$&\mathcal{F}$ Accuracy
SAM [23] 50.5 40.5
LLaVA-7B [33] 52.6 43.1
InternViT  [6] 53.9 47.5
InternVideo2 [60] 55.4 49.9
Table 5: Key component designs. Three key modules significantly boost the baseline performance (1.7, 3.6, and 2.2) separately.
ID Key Segment Context Hier. Temp. VidReasonSeg Refer-VOS
Extractor Synthesizer Synchronizer $\mathcal{J}$$&\mathcal{F}$ Accuracy $\mathcal{J}$$&\mathcal{F}$
1 48.5 39.7 59.1
2 50.2 43.8 60.7
3 52.1 43.3 63.9
4 50.7 40.9 61.5
5 53.7 48.6 65.5
6 52.2 44.8 63.0
7 54.0 46.0 66.2
8 55.4 49.9 67.5
Table 6: Hierarchical Temporal Synchronizer. Multi-scale tokens and the corresponding synchronization provide significant improvements to the baseline method. Noted that synchronization is based on multi-scale tokens.
ID Multi-scale Segment VidReasonSeg Refer-VOS
Tokens Synchronization $\mathcal{J}$$&\mathcal{F}$ Accuracy $\mathcal{J}$$&\mathcal{F}$
1 53.7 48.6 65.5
2 54.2 48.9 65.8
3 55.4 49.9 67.5
Table 7: Hierarchical Temporal Synchronizer. Adding aggregation layers will contribute to gains, but would diminish gradually.
Architecture VidReasonSeg Refer-VOS
$\mathcal{J}$$&\mathcal{F}$ Accuracy $\mathcal{J}$$&\mathcal{F}$
Baseline 53.7 48.6 65.5
+ 2 layer frame-video aggregation 54.8 49.3 66.1
+ 3 layer frame-video aggregation 55.4 49.9 67.5
+ 4 layer frame-video aggregation 55.1 49.4 67.2
@@ 4.1 Implementation Details
We use the pre-trained multimodal model from InternVideo2-1B and 6B  [60], with LoRA for fine-tuning. For the visual encoder and LLM, we apply pre-trained UMT-L from the InternVideo2 stage1 model as the vision encoder and Vicuna-7B v0. Following BLIP2, we deploy QFormer using the pre-trained $BERT_{base}$. To generate multi-scale visual features, we adopt the ViT-Adapter to generate the necessary multi-scale features for the hierarchical temporal synchronizer. Note that video segmentation differs from image segmentation in that video segmentation involves capturing inter-frame relationships across frames during training. This added complexity results in higher GPU memory requirements compared to image segmentation. To address this challenge, an efficient version of the VIT-Adapter is applied by removing all injectors.
@@
We use 8 A100 GPUs for training 50 epochs (approximately 1.5 days). We use the AdamW optimizer with the learning rate and weight decay set to $52e-$ 5 and $0.02$, respectively. The batch size per device is set to 4, the input frame number is 8, the resolution is 224, and the gradient accumulation step is set to 10. Further details are presented in the appendix.
4.2 Benchmarks
@@
Benchmarks. We evaluate ViLLa on benchmarks with video segmentation and question-answering evaluations: VideoReasonSeg, referring video segmentation (Refer-Youtube-VOS, Ref-DAVIS17), ReVOS, and conventional video segmentation (Youtube-VIS series, in the appendix). Through this evaluation, we validate the versatility of ViLLa in diverse segmentation tasks. For the VideoReasonSeg and referring video segmentation benchmark, we use the $\mathcal{J}\&\mathcal{F}$ as well as average accuracy to evaluate the video reasoning of models. In video segmentation (shown in the Appendix), we formulate the queries following the dataset’s annotations in the format of “Please segment the <description>”, with <description> corresponding to the descriptions of the instance categories. Training involves random sampling from VideoReasonSeg, VQA data, Video instance segmentation, and Referring video segmentation datasets. The model follows the training protocols of PixelLM [46], pre-training on image datasets for segmentation token initialization.
@@ 4.3 Main Results
Video Reasoning Segmentation. Tab. 1 compares ViLLa with adapted competing methods on video reasoning segmentation. For the image reasoning segmentation method, to adapt to the video segmentation, we add an extra tracker based on SAM-Track [10]. Since the original LISA only has one segmentation token, we calculate the $\mathcal{J}\&\mathcal{F}$ between the predicted tracklet and the ground truth.
@@
As shown in Tab. 1, our method outperforms the previous image reasoning segmentation LISA by a large margin, demonstrating that our model is capable of accomplishing tasks involving complex reasoning across time. Because video reasoning, different from referring video segmentation and image reasoning segmentation, requires the model to possess reasoning ability based on the whole video. Thus, without comprehensive knowledge of the whole video, the model can not perform well. What’s more, the results also indicate that our method exceeds previous reasoning VOS methods (VideoLISA, VISA) significantly. We argue that this is because of their naive design of sampling strategy and segmentation tokens that fail to model complex trajectories. On the other hand, our model improves the association between video and frame via our specially designed context synthesis strategy and decoder structures, distinguishing our method from previous works.
Figure 3: Qualitative Comparisons between ViLLa and VISA. Compared to VISA, our ViLLa successfully segments the sheep, which demonstrates better discrimination, temporal consistency, and conversation quality.
Referring Video Segmentation. Tab. 2 presents the results on referring video segmentation datasets. Our ViLLa demonstrates better performances on all these datasets, surpassing VISA by 4.5, 3.9, and 4.9 in $\mathcal{J}\&\mathcal{F}$ on Ref-Youtube-VOS, Ref-DAVIS17, and MeViS dataset, respectively.
ReVOS. Tab. 3 presents the results on ReVOS. ViLLa significantly outperforms previous methods by 4.6 $\mathcal{J}\&\mathcal{F}$, which further highlights the effectiveness of our design of synthesizer, synchronizer and extractor.
4.4 Ablation Studies
We conduct extensive ablation studies to reveal the contributions of each component. Unless otherwise specified, we conduct the experiments on our VideoReasonSeg dataset.
Design Choices of Backbone. We emphasize that different vision backbones are applicable in our framework. In Tab. 5, we show that InternVideo2 performs the best in both segmentation and multiple choices, potentially because it is the state-of-the-art video foundation model.
Key Components Design. Tab. 5 demonstrates the effect of our designs based on the backbone and transformer decoder. By adopting our method, we significantly improve the baseline by extracting key information, strengthening the video-frame relationship, and fusing context into text embedding. Tested on VideoReasonSeg, our context synthesizer improves the accuracy of multiple choices by 3.6 points, indicating that synthesizing context to the text embedding and keeping the textual input contextualized is beneficial to video understanding. Also, the design of synchronizer improves the segmentation performance by 1.7 $\mathcal{J}\&\mathcal{F}$, showing that applying multi-scale tokens and synchronizing their relations are helpful for better segmentation and tracking.
Hierarchical Temporal Synchronizer. Tab. 7 shows the influence of utilizing multi-scale tokens and video-frame synchronization in hierarchical temporal synchronizer. With multi-segment-scale tokens only, the baseline model achieves a 0.5 increase $\mathcal{J}\&\mathcal{F}$, while applying the aggregation brings about an increment of 1.7 $\mathcal{J}\&\mathcal{F}$. This demonstrates that using multi-scale tokens and synchronizing them are beneficial for video segmentation tasks. In Tab. 7, we explore the effects of increasing the number of layers in the decoder. Initially adding layers will contribute to notable gains, but the gain would diminish with more decoder layers.
4.5 Qualitative Results
In Fig. 3, we provide a visual comparison with VISA. We show that ViLLa is capable of segmenting instances (sheep with a black head) in crowded scenes with heavy occlusion. On the other hand, VISA fails to track the sheep properly after occlusion with other sheep. These results demonstrate the capacity of our ViLLa to handle complicated videos.
5 Conclusion
We introduce ViLLa, an effective and efficient model that aims to reason and segment in more complex video scenarios. We propose the key segment extractor to select pivotal frames from the video sequence, utilize the context synthesizer to generate text-related visual features, and implement the hierarchical temporal synchronizer to encourage the synchronization between video-level and frame-level segmentation tokens. Also, to help us better tune and evaluate the reasoning segmentation models, we have introduced a benchmark VideoReasonSeg, which comprises 3k video samples with both short and long videos, with both segmentation and QA as evaluations. Benefiting from our exquisite designs, ViLLa demonstrates convincing reasoning video segmentation capabilities in various video scenarios and benchmarks.
Appendix
This supplementary material provides more details about the proposed ViLLa, and our proposed benchmark, VideoReasonSeg. The first part includes discussions about the design of ViLLa and its comparison with previous methods, followed by the implementation details. Then, we provide extra ablation experiments. What’s more, we include the data generation pipeline of our VideoReasonSeg dataset, and further data cases demonstrating the variety of our data sample. Finally, we include failure case analysis. The content is organized as follows:
  • Discussions of ViLLa’s differences with previous methods.
  • The implementation details of ViLLa.
  • More ablation study experiment of ViLLa.
  • The data details and the data generation pipeline of our proposed VideoReasonSeg dataset.
  • Failure case analysis of ViLLa.
Appendix A Discussions
ViLLa is a holistically designed framework for Video Reasoning Segmentation (VRS), a burgeoning task demanding spatiotemporal tracking, segmentation, and dynamic reasoning in complex, evolving scenes. While prior works (e.g., LISA, PixelLM, Mask2Former) focus on reasoning static images or tracking through short clips, their architectures inherently mismatch VRS needs for two major reasons: 1) Temporal Reasoning Gap: Static MLLMs lack mechanisms to model motion, causality, or long-term dependencies in videos, while VOS/VIS methods lacks high-level reasoning beyond tracking. 2) Integration Challenges: Simply grafting spatial reasoning modules (e.g., ViLLa†) onto temporal trackers incurs additional computational costs and suboptimal performance. ViLLa bridges these barriers via three novel, task-specific components: i) Key Segment Extractor: Identifies critical temporal segments to reduce redundancy. ii) Context Synthesizer: Fuses and condenses long-term spatial-semantic reasoning information with dynamic scene evolution. iii) Hierarchical Temporal Synchronizer: Ensures consistency across long-term dependencies and modeling of complex scenes via multi-scale aggregation with multi-level segmentation tokens. As shown in Tab. 8, ViLLa achieves +3.5/2.4 gains over ViLLa† (adapted from LISA with designs from PixelLM, Mask2Former, and curated data), proving that adaptive integration of spatiotemporal reasoning modules—not direct reuse of static models—is essential and effective for VRS. In short, our VRS-oriented design novelly addresses understudied challenges in VRS.
Table 8: Comparison of direct adaptation of former approaches.
Method VideoReasonSeg MeViS